OpenStudy (anonymous):

hi guys can you help me to solve this if: a*d=0 x*w=0 a+x=v d+w=c v*c= what ,i know that is not always cero but how can i prove this

OpenStudy (anonymous):

\[v = a+x, ~~~ c = w+d,~~~~ so~~~ v \times c \implies (a+x)(w+d)\]

OpenStudy (anonymous):

Now use the information you have, it will be easy :) Let me know if you have any more trouble.

OpenStudy (anonymous):

than you but im sorry is x*w=0, not x*d=0

OpenStudy (anonymous):

No problem, same method though :) the equation stays just foil it out and see what you get

OpenStudy (anonymous):

ok i get v*c=(a+x)(w+d)= aw+da+xw+xd= aw+xd as da and xw equals cero what should i do now

ganeshie8 (ganeshie8):

looks it wont simplify to a number

OpenStudy (anonymous):

nope